Determine the nature of solution
y^2= 3/4y + 4/5

Answer by stanbon(75887)

Determine the nature of solution
y^2= 3/4y + 4/5
-----------------
y^2-(3/4)y-(4/5) = 0
Multiply thru by 20 to get:
------
20y^2-15y-16 = 0
----
b^2-4ac = 15^2-4*20*-16 > 0
so there are two unequal Real
Number solutions.
===================
Cheers,
Stan H.